# Energy based diffusion generator for efficient sampling of Boltzmann distributions

### v1

@@ Appendix C Importance weighting for accurate time integral
$p(t)\propto g(t)^{2}\mathbb{E}_{p(\epsilon)p_{D}(x,z_{t})}\left[\mathcal{L}_{t}(x,z_{t},\epsilon;\theta)\right].$
In our experiments, we approximate the ideal proposal using the strategy introduced in (Choi et al., 2022). The implementation details are as follows: We maintain a history buffer containing values of $\mathcal{L}_{t}$ from the most recent $B_{\mathrm{buffer}}=30$ minibatches. In each iteration, the buffer provides an estimate of the expected value of $\mathcal{L}_{t}$ as a function of $t$ using a histogram with $K{\mathrm{bin}}=100$ bins over the time interval $[0,T]$. Subsequently, we obtain an estimate of the ideal $p(t)$ for each iteration. Initially, during the early stages of training, $p(t)$ is set as a uniform distribution until $B_{\mathrm{buffer}}$ minibatches are accumulated.
Appendix D Experimental details
EDG
In our experiments, we leverage the subVP SDE proposed in (Song et al., 2020) to model the diffusion process of $z_{t}$ in (2), defined as
@@ Appendix D Experimental details
where $\zeta=(\zeta_{0},\zeta_{1})$, and $\mu_{0}(\zeta_{0};\phi),\Sigma_{0}(\zeta_{0};\phi)$ are modeled by MLPs. When incorporating the GHD in the decoder, we set $K=10$ and $J=5$. However, in the absence of GHD, denoted as ”EDG w/o GHD” in Section 4, we set $K=J=0$, and Lines 2-8 of Algorithm 1 are not implemented. Our numerical experience suggests that, for EDG without GHD based decoder, this particular setting outperforms the direct use of MLPs to model $\mu(z_{0};\phi),\Sigma(z_{0};\phi)$.
Baseline
In order to illustrate the superiority of our model, we compare EDG with long mixing time HMC, VAE with trainable Gauss Encoder (e.g. $p_{E}(z|x)=\mathcal{N}(z;\mu_{\theta}(x),\Sigma_{\theta}(x))$) and Boltzmann generator (BG) with RealNVP (Dinh et al., 2016). For the VAE, it initiates from a high-dimensional latent variable $z_{0}\in R^{30}$, comprising a three-layer fully connected network, each encompassing 256 units, interconnected using the ReLU activation function. Regarding the BG, we construct it using a three-layer fully connected structure of the RealNVP, provided with 256 units for each layer and associated with the ReLU activation function.
2D energy function
We use the following model architectures to generate samples for 2D energy tasks. ‘fc $n$’ denotes a fully connected layer with $n$ neurons.
@@
Network gauss: Score network:
    fc $2\times 32$, ReLU, 2×2$\times$2 ×(fc $32\times 32$, ReLU), fc $32\times 2$.     fc $25\times 16$, ReLU,
if GHD:     fc $16\times 16$, ReLU,
@@
Reference samples for Mog2, Mog2(i), Mog6, and Mog9 in Fig. 2 are precisely drawn from the target distributions, as these distributions are mixtures of Gaussian distributions. Reference samples for Ring and Ring5 are generated using the Metropolis-Hastings algorithm, with Gaussian distributions of variances 9 and 25, respectively, serving as proposal distributions. For the HMC, we employ the use of an HMC run of a total length of 2,000 steps, initiating from a standard normal distribution. The first 1,000 steps are designated as burn-in steps followed by a subsequent 1,000 steps used for sample generation. This procedure is independently executed 500 times for the purpose of creating the visual representation.
Given the reference samples $X=\{x_{i}\}_{i=1}^{m}$ and the generated samples $Y=\{y_{i}\}_{i=1}^{m}$, the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) measures the difference between the distributions of $X$ and $Y$ as follows:
@@
where $k(\cdot,\cdot)$ denotes the kernel function to compute the inner product. RBF kernel is used in Table 1, and the bandwidth is set as the median distance between corresponding samples. We execute 10 cycles for each model to procure 5,000 desired samples, following which we evaluate the mean and standard deviation of Maximum Mean Discrepancy (MMD) in relation to 5,000 reference data.
Bayesian Logistic Regression
In all experiments, we employ the same data partition and the datasets are divided into training and test sets at a ratio of 4:1. Before training, we normalize all datasets to have zero mean and unit variance. For $d$-dimensional features, the architectures of neural networks involved in the EDG are as follows:
@@
Network gauss: Score network:

### v2

@@ Appendix D Importance weighting for accurate time integral
$p(t)\propto g^{2}(t)\mathbb{E}_{p(\epsilon)p_{D}(x,z_{t})}\left[\mathcal{L}_{t}(x,z_{t},\epsilon;\theta)\right].$
In our experiments, we approximate the ideal proposal using the strategy introduced in [49]. The implementation details are as follows: We maintain a history buffer containing values of $\mathcal{L}_{t}$ from the most recent $30$ minibatches. In each iteration, the buffer provides an estimate of the expected value of $\mathcal{L}_{t}$ as a function of $t$ using a histogram with $100$ bins over the time interval $[0,T]$. Subsequently, we obtain an estimate of the ideal $p(t)$ for each iteration. Initially, during the early stages of training, $p(t)$ is set as a uniform distribution until $B_{\mathrm{buffer}}$ minibatches are accumulated.
Appendix E Experimental details
EDG In our experiments, we leverage the subVP SDE proposed in [37] to model the diffusion process of $z_{t}$ in (2), defined as
@@ Appendix E Experimental details
where $\zeta=(\zeta_{0},\zeta_{1})$, and $\mu_{0}(\zeta_{0};\phi),\Sigma_{0}(\zeta_{0};\phi)$ are modeled by MLPs. When incorporating the GHD in the decoder, we set $K=5$ and $J=5$.
Baseline For V-HMC, we facilitate the generation of samples after a warm-up phase comprising 1,000 steps, where each step the Metropolis-Hastings acceptance probability is computed to ensure the convergence. For L2HMC, we use a three-layer fully connected network, setting the leapfrog step length to 10. For different tasks, we adaptively select varying step sizes. For PIS, we use the publicly available code on our tasks directly. Regarding BG, the RealNVP architecture consists of 3 affine blocks, where the scaling and translation functions are modeled by a three-layer fully connected network, each with 256 units and ReLU activation functions.
@@
Network $a(\zeta;\phi)$: Score network:
      fc $10\times 32$, ReLU, 2×2$\times$2 ×(fc $32\times 32$, ReLU), fc $32\times 2$.       fc $25\times 16$, ReLU,
GHD:       fc $16\times 16$, ReLU,
@@
2D energy function We use the following model architectures to generate samples for 2D energy tasks. ‘fc $n$’ denotes a fully connected layer with $n$ neurons.
Reference samples for Mog2, Mog2(i), Mog6, and Mog9 in Fig. 2 are precisely drawn from the target distributions, as these distributions are mixtures of Gaussian distributions. Reference samples for Ring and Ring5 are generated using the Metropolis-Hastings algorithm, with a large number of iterations. Gaussian distributions with variances of 9 and 25, respectively, are used as the proposal distributions. For the V-HMC, we employ the use of an HMC run of a total length of 2,000 steps, initiating from a standard normal distribution. The first 1,000 steps are designated as burn-in steps followed by a subsequent 1,000 steps used for sample generation. This procedure is independently executed 500 times for the purpose of creating the visual representation.
Given the reference samples $X=\{x_{i}\}_{i=1}^{m}$ and the generated samples $Y=\{y_{i}\}_{i=1}^{m}$, the Maximum Mean Discrepancy (MMD) [56] measures the difference between the distributions of $X$ and $Y$ as follows:
@@
where $k(\cdot,\cdot)$ denotes the kernel function to compute the inner product. RBF kernel is used in Table 1, and the bandwidth is set as the median distance between corresponding samples. We execute 20 cycles for each model to procure 5,000 desired samples, following which we evaluate the mean value of MMD in relation to 5,000 reference data. The standard deviations are all small, hence we do not report them in the table.
@@
Network $a(\zeta;\phi)$: Score network:
@@
Bayesian Logistic Regression In all experiments, we employ the same data partition and the datasets are divided into training and test sets at a ratio of 4:1. Before training, we normalize all datasets to have zero mean and unit variance. For $d$-dimensional features, the architectures of neural networks involved in the EDG are as follows:
@@
Ising model The model architecture is shown below, which is similar to that in NeuralRG [17] without stacking bijectors to form a reversible transformation. It retains the multiscale entanglement renormalization ansatz structure, while we only use one block to update the whole dimensions of the variable directly.
The model architecture of Ising model
Decoding process Encoding process
Network $a(\zeta;\phi)$: Score network:
      hierarchy network as shown in [17]       hierarchy network.
GHD:       Input: $z_{t}$, $x$, $t$, $T$
      $Q_{v},T_{v}$:
          Conv2d $2\times 10$, BatchNorm2d, ReLU, Conv2d $10\times 1$.
      $Q_{x},T_{x}$:
          Conv2d $1\times 10$, BatchNorm2d, ReLU, Conv2d $10\times 1$.
Final network $N(\mu,\Sigma)$:
      fc $d\times 10$, ReLU, fc $10\times 10$, ReLU, fc $10\times d$.
Regarding the $p_{E}(z_{0}|x,\theta^{*})$, we employ the RK45 ODE solver implemented in scipy.integrate.solve_ivp, where the parameters set identical to those in [37]. The normalizing constant $\log Z$ is approximated according to Sec. 3.3.
Appendix F Ablation study
To elucidate the function of each component in EDG, we compare the following models on the sampling task of 2D energy functions: a VAE with the Gaussian decoder and encoder, where the means and diagonal covariance matrices are both parameterized by MLPs; a VAE with a GHD-based decoder and MLP-based encoder; EDG without GHD, where the decoder is modeled by MLPs; and the full EDG model. For VAE w/o GHD, the network is composed of 3-layer MLPs, each with 32 units and ReLU activation function. Starting from $z\sim N(0,I_{d}),d=10$, the decoder generates samples $x\sim p_{D}(x;\mu(z;\phi),\Sigma(z;\phi))$. The encoder, as an independent network, follows $p_{E}(z|x)=N(z;\mu(x;\theta),\Sigma(x;\theta))$. The objective function is the KL divergence of the joint distribution of $z$ and $x$ as
$D_{\mathrm{KL}}(p_{D}(z)p_{D}(x|z;\phi)||\pi(x)p_{E}(z|x;\theta).$
For VAE w/ GHD, the decoder is consistent with the structure in Algorithm 1, and the encoder is the same as decribed above. For EDG w/o GHD, we omit the leap-frog component (refer to Line 2-8 in the algorithm) and the decoder is composed of the network $a(\zeta;\phi)$ and a final gaussian part $\mathcal{N}(\mu,\Sigma)$ (refer to Line 1, 9 in the algorithm).
The histograms of the samples are shown in Fig. 4 for visual inspection and the sampling errors is summarized in Tab. 5. It is evident that EDG with a GHD-based decoder outperforms the other methods, demonstrating the effectiveness of each component within the model.
MoG2            MoG2(i)          MoG6            MoG9             Ring             Ring5
VAE
w/o GHD
VAE
w/ GHD
EDG
w/o GHD
EDG
w/ GHD
Figure 4: Density plots for 2D energy function. We generate $500,000$ samples for each method and plot the histogram.
Table 5: The Maximum Mean Discrepancy (MMD) between 5,000 samples generated by each generator and the reference samples.
Mog2 Mog2(i) Mog6 Mog9 Ring Ring5
VAE w/o GHD $1.86$ $1.62$ $2.57$ $2.10$ $0.12$ $0.23$
VAE w/ GHD $\mathbf{0.01}$ $2.43$ $0.15$ $0.59$ $1.68$ $0.63$
EDG w/o GHD $0.06$ $1.01$ $0.04$ $0.05$ $0.02$ $0.04$
EDG $\mathbf{0.01}$ $\mathbf{0.50}$ $\mathbf{0.01}$ $\mathbf{0.02}$ $\mathbf{0.01}$ $0.02$